# The global attractor of the inelastic linear Boltzmann equation in a gravity field for Maxwell molecules

### v1

@@ 4.1 Strategy to prove the long time behaviour
For this purpose, we will consider the Fourier transform with respect to the velocity $v$ of such solutions, namely $\mathcal{F}(t,\xi)=\int_{\mathbb{R}^{d}}f(t,v)e^{-iv\cdot\xi}\mathop{}\!\mathrm{d}v$. Since the mass of the solutions is preserved along time, we can consider probability measures for all time, and therefore the Fourier transform $\mathcal{F}(t,\xi)$ of the solution $f(t,v)$ satisfies $\mathcal{F}(t,0)=1$ and
@@
since $f\geq 0$. We recall that in probability theory the Fourier transforms of probability measures are often referred to as characteristic functions, see e.g. [36]. They are a subset $\Phi$ of the space of the complex-valued continuous functions $\mathcal{C}(\mathbb{R}^{d},\mathbb{C})$. First we consider the evolution equation of the Fourier transform $\mathcal{F}$, which reads
@@
and we refer to Appendix A and Appendix B for the detailed justification of the evolution equation for the Fourier transform (4.2).
@@
  • We will prove that initial data $f_{0}$ of the original evolution equation (1.11) can be properly compared to the steady state in Fourier variables. Namely, we will show that if $f_{0}$ has finite first order moments, as it is the case for the steady state $f_{\infty}$, then, the respective Fourier transforms $\mathcal{F}_{0}$ and $\mathcal{F}_{\infty}$ can be compared as follows
@@
    $\mathcal{C}_{1}=\{\mathcal{F}\in\mathcal{C}_{0}(\mathbb{R}^{d},\mathbb{C})\,:\,\ ||\mathcal{F}||_{\infty,1}<+\infty\},$ (4.5)
@@
    $||\mathcal{F}||_{\infty,1}=\sup_{\xi\in\mathbb{R}^{d}}\frac{|\mathcal{F}(\xi)|}{1+|\xi|}\cdotp$ (4.6)
@@ 4.2 Comparison of the Fourier transforms of two functions with same mass and finite first moments
First, we will rely on the following elementary result.
@@
$\left|\sum_{n=2}^{+\infty}\frac{\left(-v\cdot\xi\right)^{n}}{n!}\right|\leq C\min\left(|v||\xi|,|v|^{2}|\xi|^{2}\right).$ (4.8)
Proof.
The proof is obtained by distinguishing the case $|v||\xi|<1$ and $|v||\xi|\geq 1$ First we assume that $|v||\xi|<1$. In such a case, we have
$\left|\sum_{n=2}^{+\infty}\frac{\left(-v\cdot\xi\right)^{n}}{n!}\right|$ $=\left|(-iv\cdot\xi)^{2}\sum_{n=0}^{+\infty}\frac{(-iv\cdot\xi)^{n}}{(n+1)(n+2)n!}\right|$
$\leq\frac{|v|^{2}|\xi|^{2}}{2}\sum_{n=0}^{+\infty}\frac{\left(|v||\xi|\right)^{n}}{n!}=\frac{|v|^{2}|\xi|^{2}}{2}e^{|v||\xi|}$
$\leq\frac{e}{2}|v|^{2}|\xi|^{2}.$ (4.9)
If now we assume that $|v|\cdot|\xi|\geq 1$, we have
$\left|\sum_{n=2}^{+\infty}\frac{\left(-v\cdot\xi\right)^{n}}{n!}\right|$ $=\left|e^{-iv\cdot\xi}-1+iv\cdot\xi\right|$
$\leq 2+|v\cdot\xi|$
$\leq 3|v|\cdot|\xi|,$ (4.10)
where the first inequality we have used that $v\cdot\xi$ is a real number so that $|e^{-iv\cdot\xi}|=1$ and where we have used that $|v||\xi|\geq 1$ in the second inequality. Taking $C=\max\left\{\frac{e}{2},3\right\}$ concludes the proof. ∎
@@
where $C$ is the constant provided by Lemma 4.1. The proof is then complete.
@@ 4.3 Well-posedness of the evolution equation of the Fourier transform
Therefore, we introduce the following definition.
@@
By induction we show that $(\mathcal{F}^{(n)})_{n\geq 0}\in X$. In fact, we observe that trivially $\mathcal{F}^{(0)}\in X$ and that, since $\mathcal{F}_{0}\in\mathcal{C}_{1}$, we have $|\mathcal{F}^{(1)}(t,\xi)|\leq|\mathcal{F}_{0}(\xi)|$, and thus $\mathcal{F}^{(1)}\in X$. Assume now that $\mathcal{F}^{(n)}\in\widetilde{\mathcal{C}}_{1}$, we have that
@@
$\leq||\mathcal{F}_{0}||_{\infty,1}+\|\mathcal{F^{(n)}}\|_{X}\int_{0}^{t}\int_{\mathbb{S}^{d-1}}b\left(\left|\frac{\xi}{|\xi|}\cdot\sigma\right|\right)\frac{1+|\overline{\xi}|}{1+|\xi|}e^{-(t-s)}\mathop{}\!\mathrm{d}\sigma\mathop{}\!\mathrm{d}s.$ (4.24)
@@
$\left|\frac{\mathcal{F}^{(n+1)}(t,\xi)}{1+|\xi|}\right|$ $\leq||\mathcal{F}_{0}||_{\infty,1}+\|\mathcal{F}^{(n)}\|_{X}(1-e^{-t})\int_{\mathbb{S}^{d-1}}b\left(\frac{\xi}{|\xi|}\cdot\sigma\right)\mathop{}\!\mathrm{d}\sigma\leq\|\mathcal{F}_{0}\|_{\infty,1}+\|\mathcal{F}^{(n)}\|_{X}.$ (4.26)
Thus, taking the supremum first for $\xi\in\mathbb{R}^{d}$ and then for $t\in[0,T]$ in the above equation implies that $\mathcal{F}^{(n+1)}\in X$. By induction, then $(\mathcal{F}^{(n)})_{n\geq 0}\in X$ for every $n\in\mathbb{N}$.
We show now that the sequence $\left(\mathcal{F}^{(n)}\right)_{n\geq 0}$ is convergent in $X$, for $T$ appropriately chosen. We first estimate the distance of two consecutive terms of $(\mathcal{F}^{(n)})_{n\geq 0}$. In fact, we have that
@@
$\leq\|\mathcal{F}^{(n+2)}-\mathcal{F}^{(n+2)}\|_{X}\int_{0}^{t}\int_{\sigma\in\mathbb{S}^{d-1}}b\left(\left|\frac{\xi}{|\xi|}\cdot\sigma\right|\right)\frac{1+|\overline{\xi}|}{1+|\xi|}e^{-(t-s)}\mathop{}\!\mathrm{d}\sigma\mathop{}\!\mathrm{d}s$
$\leq(1-e^{-t})|\|\mathcal{F}^{(n+2)}-\mathcal{F}^{(n+2)}\|_{X}\leq(1-e^{-T})\|\mathcal{F}^{(n+2)}-\mathcal{F}^{(n+2)}\|_{X},$ (4.27)
@@
which proves that $\left(\mathcal{F}^{(n)}\right)_{n\geq 0}$ is Cauchy in $X$ and thus it is converging to a certain $\mathcal{F}\in X$. Let us observe that the chosen $T$ does not depend on $\mathcal{F}_{0}T$ therefore, we can bootstrap the previous argument, and construct recursively a solution on $[nT,(n+1)T]$, for all $n\in\mathbb{N}$. Moreover we notice that for $T=\log 2$, (4.26) implies that
@@
from which we immediately deduce by recursion
$\|\mathcal{F}^{(n)}\|_{X}\leq\|\mathcal{F}_{0}\|_{\infty,1}\sum_{k=0}^{n}\frac{1}{2^{k}}=2\|\mathcal{F}_{0}\|_{\infty,1}.$ (4.32)
@@
This is a Grönwall-type inequality, that we can study by noticing that
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\left[\left(\int_{0}^{t}||\mathcal{F}(s,\cdotp)||_{\infty,1}e^{s}\mathop{}\!\mathrm{d}s\right)e^{-t}\right]=||\mathcal{F}(t,\cdotp)||_{\infty,1}-\left(\int_{0}^{t}||\mathcal{F}(s,\cdotp)||_{\infty,1}e^{s}\mathop{}\!\mathrm{d}s\right)e^{-t}\leq||\mathcal{F}_{0}||_{\infty,1}e^{-t},$ (4.34)
from which integrating in time
$\left(\int_{0}^{t}||\mathcal{F}(s,\cdotp)||_{\infty,1}e^{s}\mathop{}\!\mathrm{d}s\right)e^{-t}\leq||\mathcal{F}_{0}||_{\infty,1}\int_{0}^{t}e^{-s}\mathop{}\!\mathrm{d}s=||\mathcal{F}_{0}||_{\infty,1}\left(1-e^{-t}\right).$ (4.35)
@@
$||\mathcal{F}(t,\cdotp)||_{\infty,1}\leq||\mathcal{F}_{0}||_{\infty,1}.$ (4.36)
We turn now to the question of uniqueness. Let $\mathcal{F},\mathcal{G}\in X$ be two solutions of (4.16) with the same initial datum $\mathcal{F}_{0}$. By definition we have
$\mathcal{F}(t,\xi)-\mathcal{G}(t,\xi)=\int_{0}^{t}\int_{\mathbb{S}^{d-1}}b\left(\left|\frac{\xi}{|\xi|}\cdot\sigma\right|\right)\left[\mathcal{F}(s,\overline{\xi})-\mathcal{G}(s,\overline{\xi})\right]e^{-(1+i(a\cdot\xi))(t-s)}\mathop{}\!\mathrm{d}\sigma\mathop{}\!\mathrm{d}s.$ (4.37)
Therefore, we have
$||\mathcal{F}(t,\cdotp)-\mathcal{G}(t,\cdotp)||_{\infty,1}\leq\int_{0}^{t}||\mathcal{F}(s,\cdotp)-\mathcal{G}(s,\cdotp)||_{\infty,1}e^{-(t-s)}\mathop{}\!\mathrm{d}s.$ (4.38)
This is another Grönwall-type inequality, that can be written as
$\left(||\mathcal{F}(t,\cdotp)-\mathcal{G}(t,\cdotp)||_{\infty,1}e^{t}\right)e^{-t}-\left(\int_{0}^{t}||\mathcal{F}(s,\cdotp)-\mathcal{G}(s,\cdotp)||_{\infty,1}e^{s}\mathop{}\!\mathrm{d}s\right)e^{-t}$
$\hskip 142.26378pt=\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\left[\left(\int_{0}^{t}||\mathcal{F}(s,\cdotp)-\mathcal{G}(s,\cdotp)||_{\infty,1}e^{s}\mathop{}\!\mathrm{d}s\right)e^{-t}\right]\leq 0.$ (4.39)
Therefore we have that the function ${t\mapsto\left(\int_{0}^{t}||\mathcal{F}(s,\cdotp)-\mathcal{G}(s,\cdotp)||_{\infty,1}e^{s}\mathop{}\!\mathrm{d}s\right)e^{-t}}$ is non-negative, decreasing, and initially zero, hence it is identically zero, as well as $t\mapsto||\mathcal{F}(t,\cdotp)-\mathcal{G}(t,\cdotp)||_{\infty,1}$. This proves the uniqueness of the solution in $X$ for arbitrary $T>0$, concluding the proof. ∎
@@
Since we are considering Fourier transforms of the solutions $f$ of (1.11), we are working with complex-valued functions in general. Nevertheless, there is an additional difficulty with respect to the situation studied in [17]: the drift term in the evolution equation (1.11) forces $f$ to be complex-valued, in the sense that, even if we assume that the initial datum $\mathcal{F}_{0}$ is real-valued, the associated solution of (4.16) will not be real-valued in general. This difficulty has to be overcome in order to define a notion of super-solution. In most cases (as it is the case in [17]), the solutions of evolutionary PDEs remain real-valued if the initial datum is real-valued. This leads to natural definitions of super-solutions, and which enables later to compare the solutions using the monotonocity of appropriate integral operators. In the present situation, it seems that there is no clear way to consider real-valued solutions of (4.16), neither to consider transformations of such solutions (such as multiplying the solutions by $e^{i(a\cdot\xi)t}$) that remain real-valued. As a consequence, we will be lead to consider a weaker notion of super-solutions. The notion we will introduce is weaker in the sense that a solution of (4.16) will not be a super-solution in the sense that will be introduced below (although our notion of super-solution requires to fulfill a more demanding inequality as the classical notion of super-solution). Another consequence of the impossibility to consider real-valued solutions is that the existence of a principle of comparison of two solutions, starting from two ordered initial data, seems to be lacking here. Nevertheless, we will be able to prove a comparison principle between a solution and a super-solution, assuming that the solution starts initially below the super-solution.
@@
We now prove the principle of comparison between a solution of (4.16), and a super-solution in the sense of Definiton 6, assuming that the initial data are ordered.
Proposition 4.4 (Principle of comparison).
@@
Consider the solution $\mathcal{F}$ to (4.2) obtained in Proposition 4.3 and assume that there exists a super-solution to the evolution equation (4.16) with initial datum $\mathcal{G}_{0}$.
Then, for all $t\geq 0$ and $\xi\in\mathbb{R}^{d}$, we have
@@
We consider again the sequence of functions $\left(\mathcal{F}^{(n)}\right)_{n\geq 0}$ introduced in (4.21), defined recursively as
@@
According to the proof of Proposition 4.3, we know that $\left(\mathcal{F}^{(n)}\right)_{n\geq 0}$ converges towards the unique solution $\mathcal{F}$ of (4.16) with initial datum $\mathcal{F}_{0}$. Since $\mathcal{G}$ is assumed to be non-negative, we have by assumption
$\left|\mathcal{F}^{(0)}(t,\xi)\right|\leq\mathcal{G}(t,\xi),$ (4.46)
for all $t\geq 0$ and $\xi\in\mathbb{R}^{d}$. Now assume that for a certain index $n\in\mathbb{N}$ we have that
@@
$\leq\mathcal{G}_{0}(\xi)e^{-t}+\int_{0}^{t}\int_{\sigma\in\mathbb{S}^{d-1}}b\left(\left|\frac{\xi}{|\xi|}\cdot\sigma\right|\right)\mathcal{G}(s,\overline{\xi})e^{-(t-s)}\mathop{}\!\mathrm{d}\sigma\mathop{}\!\mathrm{d}s$
$\leq\mathcal{G}(t,\xi),$ (4.48)
where in the last line we have used the assumption that $\mathcal{G}$ is a super-solution. Passing to the limit $n\rightarrow+\infty$ in the last inequality, we obtain that the solution $\mathcal{F}$ satisfies

### v2

@@ 4.1 Strategy to prove the long time behaviour
For this purpose, we will consider the Fourier transform with respect to the velocity $v$ of such solutions, defined as $\mathcal{F}(t,\xi)=\int_{\mathbb{R}^{d}}f(t,v)e^{-iv\cdot\xi}\mathop{}\!\mathrm{d}v$. Since the mass of the solutions is preserved along time, we can consider probability measures for all time, and therefore the Fourier transform $\mathcal{F}(t,\xi)$ of the solution $f(t,v)$ satisfies $\mathcal{F}(t,0)=1$ for all times $t$ and
@@
since $f\geq 0$. We recall that in probability theory, the Fourier transforms of probability measures are often referred to as characteristic functions, see e.g. [30]. The characteristic functions form a subset $\Phi$ of the space of the complex-valued continuous functions $\mathcal{C}(\mathbb{R}^{d},\mathbb{C})$. First we consider the evolution equation of the Fourier transform $\mathcal{F}$, which reads
@@
We refer to [21] for more details on the expression of $\overline{\xi}$, and to Appendix A for the detailed justification of the evolution equation for the Fourier transform (4.2).
@@
  • We will prove that initial data $f_{0}$ of the original evolution equation (1.9) can be properly compared to the steady state in Fourier variables. Namely, we will show that if $f_{0}$ has finite first order moments, as it is the case for the steady state $f_{\infty}$, then, their respective Fourier transforms $\mathcal{F}_{0}$ and $\mathcal{F}_{\infty}$ can be compared as follows
@@
[…]
@@
    $|\hskip-1.42262pt|\mathcal{F}|\hskip-1.42262pt|_{\infty,1}=\sup_{\xi\in\mathbb{R}^{d}}\frac{|\mathcal{F}(\xi)|}{1+|\xi|}\cdotp$ (4.6)
@@ 4.2 Comparison of the Fourier transforms of two functions with same mass and finite first moments
We will rely on the following result.
@@
$\left|\sum_{n=2}^{+\infty}\frac{\left(-iv\cdot\xi\right)^{n}}{n!}\right|\leq C\min\left(|v||\xi|,|v|^{2}|\xi|^{2}\right).$ (4.8)
The proof, which is elementary, is obtained by distinguishing the cases $|v||\xi|<1$ and $|v||\xi|\geq 1$.
@@
where $C$ is the constant provided by Lemma 4.1. The proof is then complete. ∎
@@ 4.3 Well-posedness of the evolution equation of the Fourier transform
We introduce the following definition.
@@
Remark 5.
Observe that the result of Proposition 4.3 provides the well-posedness result for the inelastic linear Boltzmann equation (1.9), via a different approach from the semigroup strategy of Section 2. The drawbacks of this approach are, on the ond hand more stringent hypotheses, and on the other hand, the fact that one needs to prove that the solution that is obtained for (4.2) corresponds indeed to the Fourier transform of a probability measure, as it is carefully done in [15].
@@
By induction we show that $(\mathcal{F}^{(n)})_{n\geq 0}\in X$. In fact, we observe that trivially $\mathcal{F}^{(0)}\in X$ and that, since $\mathcal{F}_{0}\in\mathcal{C}_{1}$, we have $|\mathcal{F}^{(1)}(t,\xi)|\leq|\mathcal{F}_{0}(\xi)|$, and thus $\mathcal{F}^{(1)}\in X$. Assume now that $\mathcal{F}^{(n)}\in X$, we have that
@@
[…]
@@
[…]
Thus, taking the supremum first on $\xi\in\mathbb{R}^{d}$ and then on $t\in[0,T]$ in the above equation implies that $\mathcal{F}^{(n+1)}\in X$. By induction, we deduce that $(\mathcal{F}^{(n)})_{n\geq 0}\in X$ for every $n\in\mathbb{N}$.
We show now that the sequence $\left(\mathcal{F}^{(n)}\right)_{n\geq 0}$ is converging in $X$, for $T$ appropriately chosen. We first estimate the distance of two consecutive terms of $(\mathcal{F}^{(n)})_{n\geq 0}$. In fact, we have that
@@
$\leq\|\mathcal{F}^{(n+1)}-\mathcal{F}^{(n)}\|_{X}\int_{0}^{t}\int_{\sigma\in\mathbb{S}^{d-1}}b\left(\left|\frac{\xi}{|\xi|}\cdot\sigma\right|\right)\frac{1+|\overline{\xi}|}{1+|\xi|}e^{-(t-s)}\mathop{}\!\mathrm{d}\sigma\mathop{}\!\mathrm{d}s$
$\leq(1-e^{-t})|\hskip-1.42262pt\|\mathcal{F}^{(n+1)}-\mathcal{F}^{(n)}\|_{X}\leq(1-e^{-T})\|\mathcal{F}^{(n+1)}-\mathcal{F}^{(n)}\|_{X},$ (4.25)
@@
which proves that $\left(\mathcal{F}^{(n)}\right)_{n\geq 0}$ is a Cauchy sequence in $X$ and thus it is converging to a certain $\mathcal{F}\in X$. Let us observe that the chosen $T$ does not depend on $\mathcal{F}_{0}$ and therefore, we can bootstrap the previous argument, and construct recursively a solution on $[nT,(n+1)T]$, for all $n\in\mathbb{N}$. Moreover we notice that for $T=\log 2$, (4.24) implies that
@@
from which we deduce by an immediate recursion
$\|\mathcal{F}^{(n)}\|_{X}\leq\sum_{k=0}^{n}\frac{1}{2^{k}}\|\mathcal{F}_{0}\|_{\infty,1}=2\|\mathcal{F}_{0}\|_{\infty,1}.$ (4.29)
@@
This is a Grönwall-type inequality, which can be rewritten as
[…]
from which integrating in time provides
[…]
@@
[…]
We turn now to the question of the uniqueness. Let $\mathcal{F},\mathcal{G}\in X$ be two solutions of (4.14) with the same initial datum $\mathcal{F}_{0}$. The equation (4.14) being linear, $\mathcal{F}-\mathcal{G}$ is also a solution to (4.14), with identically zero initial datum. Applying (4.33) to $\mathcal{F}-\mathcal{G}$, we deduce that the difference $\mathcal{F}-\mathcal{G}$ is zero for any time. This proves the uniqueness of the solution in $X$ for arbitrary $T>0$, concluding the proof. ∎
@@
Since we are considering Fourier transforms of the solutions $f$ of (1.9), we are working with complex-valued functions in general. In most cases (as it is the case in [15]), the solutions of evolutionary PDEs remain real-valued if the initial datum is real-valued. This leads to natural definitions of super-solutions, and which enables later to compare the solutions using the monotonocity of appropriate integral operators.
In the present case, there is an additional difficulty with respect to the situation studied in [15]: the drift term in the evolution equation (1.9) forces $f$ to be complex-valued, in the sense that, even if we assume that the initial datum $\mathcal{F}_{0}$ is real-valued, the associated solution of (4.14) will not be real-valued in general. Moreover, it seems that there is no clear way to consider real-valued solutions of (4.14), neither to consider transformations of such solutions that would remain real-valued (such as multiplying the solutions by $e^{i(a\cdot\xi)t}$).
As a consequence, we will be led to introduce a weaker notion of super-solution. This notion is weaker in the sense that a solution of (4.14) will not be a super-solution under the forthcoming definition, although our notion involves a more demanding inequality than the classical one. Moreover, since real-valued solutions cannot be considered, no comparison principle between two ordered solutions seems to hold. Nevertheless, we will establish a comparison principle between a solution and a super-solution, provided that the former initially lies below the latter.
@@
We now prove the comparison principle between a solution of (4.14), and a super-solution in the sense of Definiton 6, assuming that the initial data are ordered.
Proposition 4.4 (Comparison principle).
@@
Consider the solution $\mathcal{F}$ to (4.2) obtained in Proposition 4.3 with initial datum $\mathcal{F}_{0}$ and assume that there exists a super-solution to (4.14) in the sense of Definition 6, with initial datum $\mathcal{G}_{0}$. Then, for all $t\geq 0$ and $\xi\in\mathbb{R}^{d}$, we have
@@
We consider again the sequence of functions introduced in (4.19), defined recursively as
@@
According to the proof of Proposition 4.3, we know that $\left(\mathcal{F}^{(n)}\right)_{n\geq 0}$ converges towards the unique solution $\mathcal{F}$ of (4.14) with initial datum $\mathcal{F}_{0}$. Since $\mathcal{G}$ is assumed to be non-negative, we have by assumption $\left|\mathcal{F}^{(0)}(t,\xi)\right|\leq\mathcal{G}(t,\xi)$ for all $t\geq 0$ and $\xi\in\mathbb{R}^{d}$. Now assume that for a certain index $n\in\mathbb{N}$ we have that
@@
$\leq\mathcal{G}_{0}(\xi)e^{-t}+\int_{0}^{t}\int_{\sigma\in\mathbb{S}^{d-1}}b\left(\left|\frac{\xi}{|\xi|}\cdot\sigma\right|\right)\mathcal{G}(s,\overline{\xi})e^{-(t-s)}\mathop{}\!\mathrm{d}\sigma\mathop{}\!\mathrm{d}s\leq\mathcal{G}(t,\xi),$ (4.41)
where in the last line we have used the assumption that $\mathcal{G}$ is a super-solution. We deduce by recursion that (4.40) holds for any index $n\in\mathbb{N}$. Passing to the limit $n\rightarrow+\infty$ in the (4.40), we obtain that the solution $\mathcal{F}$ satisfies